# Identifying Dwarfs of MC Analog GalaxiEs (ID-MAGE): The Search for Satellites Around Low-Mass Hosts

### v3

@@ 5.2 Sample Contamination
Contamination is a well-documented challenge for unresolved satellite galaxy searches (e.g., Bennet et al. 2019, 2020; Carlsten et al. 2022). In some cases, like NGC 3738 and NGC 4244, likely background galaxy groups can be identified within the candidate sample. These hosts exhibit an unusually high concentration of satellite candidates in small regions of their search area. Many candidates around NGC 3738 are likely associated with NGC 3718, a Seyfert 1 galaxy at 14 Mpc, or a galaxy cluster at 21 Mpc including NGC 3733 and NGC 3756, based on their close proximity to the massive galaxies and each other. Due to its potentially higher contamination rate, we exclude NGC 3738 from further analysis to prevent skewing our sample results.
@@
While the sample certainly includes some false positives, there is strong evidence that a significant portion are real satellites. By spatially binning the satellite sample into equal-area annular bins, we observe a distinct concentration of candidates within 60 kpc of their host, compared to the outer bins (see Figure 7). The number of satellites per bin should correlate with the distance from the host. However, the number of background galaxies should be roughly constant as the number should only depend on area and the annular rings are equal in area. This assumption of a uniform surface density of the background contamination has been used and tested previously for imaging based satellite searches around galaxies with masses of 9.5log(M∗/M⊙)11 (e.g., Wu et al. 2022). However, the background distribution for LMC-mass systems may not be as uniform compared higher-mass systems. The central concentration seen in Figure 7 aligns with the models from Dooley et al. (2017b, a), which are based on the Caterpillar simulation suite (Griffen et al., 2016). They predict that approximately 50–65% of satellites around an LMC-mass galaxy should lie within 50–60 kpc of their host. This central concentration is even more pronounced in the high-likelihood sample, suggesting it has a lower contamination rate. The ELVES and SAGA satellite samples demonstrate a clear central concentration in Figure 9 of Carlsten et al. (2022) which compares the number of satellites per host in equal width annular bins. For the satellites in the ELVES sample, the central concentration is strongest for the more massive satellites, which is not seen for the SAGA satellite sample.
@@
The number of background contaminants in each bin should be roughly equal. To estimate the number of background galaxies per bin, we use the outer bins in Figure 7. For the high-likelihood sample, the average number of contaminants per host per bin in the outer five bins for our LMC-mass hosts is 0.490.10 for LMC-mass hosts and 0.420.06 for our SMC-mass hosts. The number for the innermost bin is 1.40.40 for LMC-mass hosts and 1.120.21 for SMC-mass hosts. After accounting for the average number of contaminants from the outer bins, the innermost bin shows an excess of 0.910.41 candidates for LMC-mass hosts and 0.700.22 candidates for SMC-mass hosts. If we assume that approximately 60% of our satellites reside within 60 kpc of their host (Dooley et al., 2017a), this provides a lower estimate of 1.60.7 (1.20.4 ) identified satellites per LMC-mass (SMC-mass) host in the high-likelihood sample. This aligns well with the lower range of cosmological predictions from Dooley et al. (2017b) (see Section 5.3), and the lower estimate for our LMC-mass hosts is in agreement with the satellite luminosity function of NGC 2403 (an LMC-mass host) in Carlin et al. (2024).
5.3 Satellite Luminosity Function
@@
Satellite surveys have found a correlation between the satellite abundance and stellar mass of the host for MW-mass galaxies (Mao et al., 2021; Carlsten et al., 2021; Danieli et al., 2023; Mutlu-Pakdil et al., 2024). Hydrodynamical simulations using different abundance matching models and stellar-mass to halo-mass relations () broadly predict that as the host’s halo mass increases, the stellar mass increases and the number of satellites within a given mass range also increases (e.g., Santos-Santos et al. 2022). Carlsten et al. (2022) found that the simulated luminosity functions from the ARTEMIS simulations (Font et al., 2021) agree well with the average satellite abundance of MW-mass hosts at higher mass. However, the predictions may modestly underestimate the satellite abundance of the less massive hosts. This demonstrates a need to constrain the satellite abundances of galaxies less massive than the MW. Observations of the satellite luminosity functions (LF) of dwarf galaxies provide vital tests to different abundance-matching models and constraints to the slope of the relationship.
@@
To estimate an upper range of our hosts’ LF, we consider a maximum number of satellites we detect. We assume that the visual inspection is entirely reliable and that all high-likelihood candidates will be confirmed as satellites. We combine the high-likelihood sample with the already confirmed satellites as our upper estimate. As discussed in Section 5.2, our high-likelihood sample consists of the candidates with lower contamination from background and foreground galaxies. Additionally, for SMC-mass hosts, we apply a virial radius cutoff of 110 kpc —slightly larger than the estimated virial radius of the SMC (Mutlu-Pakdil et al., 2021)—to exclude probable interlopers as the area searched extends to 150 kpc. Figure 8 presents the LF for our hosts with these assumptions. To account for variations in satellite abundance between hosts, the error bars are the standard deviation of the number of satellites per magnitude bin per host, divided by . Figure 8 also shows the LF based on our lower estimates of the number of satellites per host from the central concentration plots.
@@
For the LMC and SMC-mass hosts, the LF appears to rise as the absolute magnitude becomes fainter. Overall, LMC-mass hosts tend to have more satellites per magnitude bin than SMC-mass hosts. This is consistent with simulations (Dooley et al., 2017a, b; Santos-Santos et al., 2022) which predict LMC-mass galaxies host richer satellite systems than SMC-mass galaxies. This difference is not due to the larger virial radius used for LMC-mass hosts. Expanding the SMC-mass host radius to 150 kpc slightly increases the number of satellite candidates per bin, but the overall trend remains unchanged.
@@
Figure 8 illustrates that our hosts have significantly fewer satellites per magnitude bin than the MW-mass hosts of the ELVES survey. Comparing the three host mass ranges, there is a clear trend that as the host mass increases, so does the number of satellites in each magnitude bin. This trend aligns well with predictions from the CDM model that satellite abundance correlates with host mass (e.g., Dooley et al. 2017b, a; Santos-Santos et al. 2022). This also agrees with the correlation Carlsten et al. (2021) observed between the stellar masses of MW-mass hosts and their satellite abundances. Our sample extends this observed trend to the dwarf host galaxy mass range.
@@
In Figure 9, we compare our LF with theoretical predictions taken from Dooley et al. (2017a). These predictions are derived from the dark matter-only Caterpillar simulation suite (Griffen et al., 2016), combined with different - relations (Moster et al., 2013; Brook et al., 2014; Garrison-Kimmel et al., 2017). Since most of our satellites are unconfirmed, our LF represents an upper limit of the true luminosity function. On average, LMC-mass hosts contain 4.01.4 satellites per host, while SMC-mass hosts have 2.10.6 satellites per host with M-. The shaded region shows the standard deviation of the number of satellites per magnitude bin per host, divided by .
@@
The upper estimate of our LF trends toward the higher end of predictions, particularly for the LMC-mass hosts. The upper-estimate LF falls within the 1 range of predictions based on the Garrison-Kimmel et al. (2017) relation. The more massive satellites–where a larger fraction are confirmed –show good agreement with predictions for both LMC- and SMC-mass hosts. Further, our upper estimate for the LF of the SMC-mass hosts is in agreement with the LF of NGC 3109 (a SMC-mass host) in Doliva-Dolinsky et al. (2025) Our lower limits of 1.60.7 (1.20.4 ) candidates per LMC-mass (SMC-mass) host with M- are similar to the number satellite identified and confirmed per host by ELVES-Dwarf (Li et al., 2025). Li et al. (2025) confirmed zero to two satellites per SMC-mass host and zero to one satellites per LMC-mass host.
@@
Our upper-estimate is an over-estimate; however, more than half of the high-likelihood candidates would need to be interlopers for the final confirmed sample to be below the lower range of the predictions in Dooley et al. (2017a). As seen in Figure 9, our lower estimates from Section 5.2, 1.20.4 (SMC-mass) and 1.60.7 (LMC-mass), align with the lower-range of predictions from Dooley et al. (2017a) and Dooley et al. (2017b). Therefore, the number of confirmed satellites after our follow-up observations is expected to fall within these bounds. Additionally, the true shape of the LF is likely different due to the interlopers in our sample and because we have not applied a completeness correction. The flattening of the LF for the lowest mass bin is likely due to the lower completeness in that bin, which can be attributed in part to the lower detection completeness for our more distant hosts. Once finalized, our sample will enable a more thorough comparison between observations and theoretical predictions, offering the first statistically significant observational constraints on the number of satellites around LMC- and SMC-mass hosts.
6 Follow-Up Campaign
@@
The ultimate goal of ID-MAGE is to create the first statistical view of dwarf satellites around low-mass hosts, substantially extending the range of host masses and environments probed by existing surveys (see Figure 1), delivering quantitative constraints for galaxy formation physics in CDM. We are currently conducting a comprehensive observational follow-up campaign to confirm and characterize our satellite candidates. As a first step, we check our candidates against the H i surveys HIPASS (Meyer et al., 2004) and ALFALFA (Haynes et al., 2018) because if we can identify H i line emission, we can immediately confirm or refute the association with the host. For those without H i redshifts, we are leading two complementary follow-up campaigns to measure their distances: (1) deeper imaging follow-up for the SBF technique, and (2) optical spectroscopic follow-up for redshift measurements. This will allow us to confidently identify and remove interlopers, resulting in a clean catalog of satellites around our host galaxies. To characterize our satellites, we are leading an H i observation campaign, coupled with a GALEX UV photometric study, to study their gas content and star formation activity. Additionally, we will compare satellite and system properties (e.g., star formation rates, quenched fraction, gas fraction) with recent studies investigating the environmental impact of MW-host on dwarf galaxies (e.g., Karunakaran et al., 2020, 2021, 2022, 2023; Jones et al., 2024). Our follow-up campaign will enable us to examine the gas retention and star formation activity in individual satellites, as well as the quenched fraction and luminosity function across the satellite systems.
@@
While visual inspection effectively reduces false positives, contamination from background galaxies often remains significant, and can account for a large proportion of satellite candidates (e.g., Carlsten et al. 2022; Bennet et al. 2019, 2020). The contamination rate can be as high as 80% in searches around MW-mass galaxies in cases where there is a massive galaxy in the background with its own satellite system. Thus, follow-up observations are essential to verify that these candidates are genuine satellites of the purported hosts. Our follow-up campaign is actively underway, with several current and recently completed observational runs.
@@
SBF offers an efficient way to measure distances to quenched dwarfs (i.e., ones without star-forming regions) in the distance range of our hosts (Carlsten et al., 2019a, b). Although the DESI Legacy Imaging Surveys provide full virial volume coverage of our host galaxies, which is essential for a systemic dwarf search, the data are too shallow and the seeing is too poor to perform SBF distance measurements. Therefore, we are leading a deep follow-up imaging campaign with MMT, Gemini North, and Magellan to use the SBF technique efficiently.
@@
It is well known that the SBF technique is not ideal for gas-rich, star-forming systems because their star-forming regions can significantly affect the SBF measurements (Greco et al., 2021). Therefore, for our blue candidates with patchy morphology, we are simultaneously undergoing a ground-based spectroscopic campaign to obtain redshift measurements. Similar campaigns have successfully obtained redshifts of diffuse dwarf galaxies from clear emission (e.g., Greco et al., 2018) or absorption lines (e.g., van Dokkum et al., 2015b; Kadowaki et al., 2017). As part of this campaign, we have secured observational time to conduct long-slit spectroscopy using several telescopes, including Gemini/GMOS, LBT/MODS, Magellan/IMACS, MMT/Binospec, SALT/RSS, and MDM/OSMOS.
@@
H i is the initial fuel for star formation, so its presence or lack thereof in satellites enables a better understanding of their past and future evolution. We have been awarded 200 hours on the Green Bank Telescope (GBT; PI: Hunter) to assess whether our candidates are gas-poor, old stellar systems, or gas-rich, recently star-forming dwarfs. Our observational strategy is designed to give a near uniform sensitivity across all targets in terms of their H i -to-stellar mass ratio, with our targets being (at ). Given that for gas-rich field dwarfs (Huang et al., 2012), even for undetected targets, we will be able to confidently conclude that they are gas-poor. Furthermore, we will compare the H i measurements with star formation rates derived from archival GALEX UV data. This combined data set will enable us to determine the quenched fraction of satellites around low-mass hosts and explore the mechanisms responsible for quenching.
7 Conclusions
@@
We employ the detection algorithm described in Bennet et al. (2017) with additional visual inspection by experts. We assess the completeness of our algorithm using artificial dwarf injections. Additionally, we evaluate the completeness of our visual inspection. Combined, the algorithm and visual inspection are complete down to and mag arcsec-2.
In total, we identify 355 satellite candidates, including 264 newly discovered galaxies. Among these, 134 are classified as high-likelihood candidates based on our systematic visual inspection. Of the 134 high-likelihood candidates, 36 are associated with LMC-mass hosts and 98 with SMC-mass hosts. The number of satellite candidates per hosts ranges from 0 to 15 high-likelihood candidates. This scatter may be driven by physical factors, such as environmental richness or intrinsic host-to-host scatter. However, some of this variation is likely due to foreground and background contaminants. The candidates in our sample range in apparent magnitudes between 15.0 20.9 and effective radii between 3.1″and 75″. The photometric properties of our sample are consistent with satellites of MW-mass galaxies and dwarf galaxies within 3 Mpc (Pace, 2024).
For our hosts, we identify fewer satellites per magnitude bin per host compared to MW-mass hosts and we find fewer candidates per magnitude bin for the SMC-mass hosts compared to the LMC-mass hosts. This agrees with CDM simulations that predict as the host’s halo mass increases, the stellar mass increases, and so to does the number of satellites (e.g. Dooley et al. 2017a, b; Santos-Santos et al. 2022). Our low-mass hosts also exhibit the trend observed among MW-mass hosts, where satellite abundance correlates with host stellar mass, extending this relationship into the dwarf host galaxy mass range.
From our candidate sample, we establish upper and lower estimates of the LF for low-mass galaxies. To determine the lower estimate, we analyze the central concentration of candidates around their hosts, estimating a rough lower bound of 1.60.7 (1.20.4 ) satellites per LMC-mass (SMC-mass) host with M9. The high-likelihood sample serves as the upper estimate, with 4.01.4 (2.10.6 ) candidates per LMC-mass (SMC-mass) host. Our upper and lower estimates bracket the predicted range for satellites of dwarf galaxies from Dooley et al. (2017a, b).

### v5

@@ V.2 Sample Contamination
Contamination is a well-documented challenge for unresolved satellite galaxy searches (e.g., Bennet19; Bennet20; ELVES). In some cases, like NGC 3738 and NGC 4244, likely background galaxy groups can be identified within the candidate sample. These hosts exhibit an unusually high concentration of satellite candidates in small regions of their search area. Many candidates around NGC 3738 are likely associated with NGC 3718, a Seyfert 1 galaxy at 14 Mpc, or a galaxy cluster at 21 Mpc including NGC 3733 and NGC 3756, based on their close proximity to the massive galaxies and each other. Due to its potentially higher contamination rate, we exclude NGC 3738 from further analysis to prevent skewing our sample results.
@@
While the sample certainly includes some false positives, there is strong evidence that a significant portion are real satellites. By spatially binning the satellite sample into equal-area annular bins, we observe a distinct concentration of candidates within 60 kpc of their host, compared to the outer bins (see Figure 7). The number of satellites per bin should correlate with the distance from the host. However, the number of background galaxies should be roughly constant as the number should only depend on area and the annular rings are equal in area. This assumption of a uniform surface density of the background contamination has been used and tested previously for imaging based satellite searches around galaxies with masses of 9.5log(M∗/M⊙)11 (e.g., xSAGA). However, the background distribution for LMC-mass systems may not be as uniform compared higher-mass systems. The central concentration seen in Figure 7 aligns with the models from Dooley17a; Dooley17, which are based on the Caterpillar simulation suite (Griffin16). They predict that approximately 50–65% of satellites around an LMC-mass galaxy should lie within 50–60 kpc of their host. This central concentration is even more pronounced in the high-likelihood sample, suggesting it has a lower contamination rate. The ELVES and SAGA satellite samples demonstrate a clear central concentration in Figure 9 of ELVES which compares the number of satellites per host in equal width annular bins. For the satellites in the ELVES sample, the central concentration is strongest for the more massive satellites, which is not seen for the SAGA satellite sample.
@@
The number of background contaminants in each bin should be roughly equal. To estimate the number of background galaxies per bin, we use the outer bins in Figure 7. For the high-likelihood sample, the average number of contaminants per host per bin in the outer five bins for our LMC-mass hosts is 0.490.10 for LMC-mass hosts and 0.420.06 for our SMC-mass hosts. The number for the innermost bin is 1.40.40 for LMC-mass hosts and 1.120.21 for SMC-mass hosts. After accounting for the average number of contaminants from the outer bins, the innermost bin shows an excess of 0.910.41 candidates for LMC-mass hosts and 0.700.22 candidates for SMC-mass hosts. If we assume that approximately 60% of our satellites reside within 60 kpc of their host (Dooley17), this provides a lower estimate of 1.60.7 (1.20.4 ) identified satellites per LMC-mass (SMC-mass) host in the high-likelihood sample. This aligns well with the lower range of cosmological predictions from Dooley17a (see Section V.3), and the lower estimate for our LMC-mass hosts is in agreement with the satellite luminosity function of NGC 2403 (an LMC-mass host) in Carlin24.
V.3 Satellite Luminosity Function
@@
Satellite surveys have found a correlation between the satellite abundance and stellar mass of the host for MW-mass galaxies (SAGAII; Carlsten21a; Danieli23; BMP24). Hydrodynamical simulations using different abundance matching models and stellar-mass to halo-mass relations () broadly predict that as the host’s halo mass increases, the stellar mass increases and the number of satellites within a given mass range also increases (e.g., Santos-Santos22). ELVES found that the simulated luminosity functions from the ARTEMIS simulations (Font2021) agree well with the average satellite abundance of MW-mass hosts at higher mass. However, the predictions may modestly underestimate the satellite abundance of the less massive hosts. This demonstrates a need to constrain the satellite abundances of galaxies less massive than the MW. Observations of the satellite luminosity functions (LF) of dwarf galaxies provide vital tests to different abundance-matching models and constraints to the slope of the relationship.
@@
To estimate an upper range of our hosts’ LF, we consider a maximum number of satellites we detect. We assume that the visual inspection is entirely reliable and that all high-likelihood candidates will be confirmed as satellites. We combine the high-likelihood sample with the already confirmed satellites as our upper estimate. As discussed in Section V.2, our high-likelihood sample consists of the candidates with lower contamination from background and foreground galaxies. Additionally, for SMC-mass hosts, we apply a virial radius cutoff of 110 kpc —slightly larger than the estimated virial radius of the SMC (BMP2021)—to exclude probable interlopers as the area searched extends to 150 kpc. Figure 8 presents the LF for our hosts with these assumptions. To account for variations in satellite abundance between hosts, the error bars are the standard deviation of the number of satellites per magnitude bin per host, divided by . Figure 8 also shows the LF based on our lower estimates of the number of satellites per host from the central concentration plots.
@@
For the LMC and SMC-mass hosts, the LF appears to rise as the absolute magnitude becomes fainter. Overall, LMC-mass hosts tend to have more satellites per magnitude bin than SMC-mass hosts. This is consistent with simulations (Dooley17; Dooley17a; Santos-Santos22) which predict LMC-mass galaxies host richer satellite systems than SMC-mass galaxies. This difference is not due to the larger virial radius used for LMC-mass hosts. Expanding the SMC-mass host radius to 150 kpc slightly increases the number of satellite candidates per bin, but the overall trend remains unchanged.
@@
Figure 8 illustrates that our hosts have significantly fewer satellites per magnitude bin than the MW-mass hosts of the ELVES survey. Comparing the three host mass ranges, there is a clear trend that as the host mass increases, so does the number of satellites in each magnitude bin. This trend aligns well with predictions from the CDM model that satellite abundance correlates with host mass (e.g., Dooley17a; Dooley17; Santos-Santos22). This also agrees with the correlation Carlsten21a observed between the stellar masses of MW-mass hosts and their satellite abundances. Our sample extends this observed trend to the dwarf host galaxy mass range.
@@
In Figure 9, we compare our LF with theoretical predictions taken from Dooley17. These predictions are derived from the dark matter-only Caterpillar simulation suite (Griffin16), combined with different - relations (Moster13; Brook14; GK16). Since most of our satellites are unconfirmed, our LF represents an upper limit of the true luminosity function. On average, LMC-mass hosts contain 4.01.4 satellites per host, while SMC-mass hosts have 2.10.6 satellites per host with M-. The shaded region shows the standard deviation of the number of satellites per magnitude bin per host, divided by .
@@
The upper estimate of our LF trends toward the higher end of predictions, particularly for the LMC-mass hosts. The upper-estimate LF falls within the 1 range of predictions based on the GK16 relation. The more massive satellites–where a larger fraction are confirmed –show good agreement with predictions for both LMC- and SMC-mass hosts. Further, our upper estimate for the LF of the SMC-mass hosts is in agreement with the LF of NGC 3109 (a SMC-mass host) in ADD3109 Our lower limits of 1.60.7 (1.20.4 ) candidates per LMC-mass (SMC-mass) host with M- are similar to the number satellite identified and confirmed per host by ELVES-Dwarf (ELVES-Dwarf). ELVES-Dwarf confirmed zero to two satellites per SMC-mass host and zero to one satellites per LMC-mass host.
@@
Our upper-estimate is an over-estimate; however, more than half of the high-likelihood candidates would need to be interlopers for the final confirmed sample to be below the lower range of the predictions in Dooley17. As seen in Figure 9, our lower estimates from Section V.2, 1.20.4 (SMC-mass) and 1.60.7 (LMC-mass), align with the lower-range of predictions from Dooley17 and Dooley17a. Therefore, the number of confirmed satellites after our follow-up observations is expected to fall within these bounds. Additionally, the true shape of the LF is likely different due to the interlopers in our sample and because we have not applied a completeness correction. The flattening of the LF for the lowest mass bin is likely due to the lower completeness in that bin, which can be attributed in part to the lower detection completeness for our more distant hosts. Once finalized, our sample will enable a more thorough comparison between observations and theoretical predictions, offering the first statistically significant observational constraints on the number of satellites around LMC- and SMC-mass hosts.
VI Follow-Up Campaign
@@
The ultimate goal of ID-MAGE is to create the first statistical view of dwarf satellites around low-mass hosts, substantially extending the range of host masses and environments probed by existing surveys (see Figure 1), delivering quantitative constraints for galaxy formation physics in CDM. We are currently conducting a comprehensive observational follow-up campaign to confirm and characterize our satellite candidates. As a first step, we check our candidates against the H i surveys HIPASS (HIPASS) and ALFALFA (ALFALFA) because if we can identify H i line emission, we can immediately confirm or refute the association with the host. For those without H i redshifts, we are leading two complementary follow-up campaigns to measure their distances: (1) deeper imaging follow-up for the SBF technique, and (2) optical spectroscopic follow-up for redshift measurements. This will allow us to confidently identify and remove interlopers, resulting in a clean catalog of satellites around our host galaxies. To characterize our satellites, we are leading an H i observation campaign, coupled with a GALEX UV photometric study, to study their gas content and star formation activity. Additionally, we will compare satellite and system properties (e.g., star formation rates, quenched fraction, gas fraction) with recent studies investigating the environmental impact of MW-host on dwarf galaxies (e.g., Karunakaran20; Karunakaran21; Karaunkaran22; Karunakaran23; Jones24). Our follow-up campaign will enable us to examine the gas retention and star formation activity in individual satellites, as well as the quenched fraction and luminosity function across the satellite systems.
@@
While visual inspection effectively reduces false positives, contamination from background galaxies often remains significant, and can account for a large proportion of satellite candidates (e.g., ELVES; Bennet19; Bennet20). The contamination rate can be as high as 80% in searches around MW-mass galaxies in cases where there is a massive galaxy in the background with its own satellite system. Thus, follow-up observations are essential to verify that these candidates are genuine satellites of the purported hosts. Our follow-up campaign is actively underway, with several current and recently completed observational runs.
@@
SBF offers an efficient way to measure distances to quenched dwarfs (i.e., ones without star-forming regions) in the distance range of our hosts (Carlsten19; carlsten19a). Although the DESI Legacy Imaging Surveys provide full virial volume coverage of our host galaxies, which is essential for a systemic dwarf search, the data are too shallow and the seeing is too poor to perform SBF distance measurements. Therefore, we are leading a deep follow-up imaging campaign with MMT, Gemini North, and Magellan to use the SBF technique efficiently.
@@
It is well known that the SBF technique is not ideal for gas-rich, star-forming systems because their star-forming regions can significantly affect the SBF measurements (Greco21). Therefore, for our blue candidates with patchy morphology, we are simultaneously undergoing a ground-based spectroscopic campaign to obtain redshift measurements. Similar campaigns have successfully obtained redshifts of diffuse dwarf galaxies from clear emission (e.g., Greco2018) or absorption lines (e.g., vanDokkum2015; Kadowaki2017). As part of this campaign, we have secured observational time to conduct long-slit spectroscopy using several telescopes, including Gemini/GMOS, LBT/MODS, Magellan/IMACS, MMT/Binospec, SALT/RSS, and MDM/OSMOS.
@@
H i is the initial fuel for star formation, so its presence or lack thereof in satellites enables a better understanding of their past and future evolution. We have been awarded 200 hours on the Green Bank Telescope (GBT; PI: Hunter) to assess whether our candidates are gas-poor, old stellar systems, or gas-rich, recently star-forming dwarfs. Our observational strategy is designed to give a near uniform sensitivity across all targets in terms of their H i -to-stellar mass ratio, with our targets being (at ). Given that for gas-rich field dwarfs (Huang12), even for undetected targets, we will be able to confidently conclude that they are gas-poor. Furthermore, we will compare the H i measurements with star formation rates derived from archival GALEX UV data. This combined data set will enable us to determine the quenched fraction of satellites around low-mass hosts and explore the mechanisms responsible for quenching.
VII Conclusions
@@
We employ the detection algorithm described in Bennet17 with additional visual inspection by experts. We assess the completeness of our algorithm using artificial dwarf injections. Additionally, we evaluate the completeness of our visual inspection. Combined, the algorithm and visual inspection are complete down to and mag arcsec-2.
In total, we identify 355 satellite candidates, including 264 newly discovered galaxies. Among these, 134 are classified as high-likelihood candidates based on our systematic visual inspection. Of the 134 high-likelihood candidates, 36 are associated with LMC-mass hosts and 98 with SMC-mass hosts. The number of satellite candidates per hosts ranges from 0 to 15 high-likelihood candidates. This scatter may be driven by physical factors, such as environmental richness or intrinsic host-to-host scatter. However, some of this variation is likely due to foreground and background contaminants. The candidates in our sample range in apparent magnitudes between 15.0 20.9 and effective radii between 3.1″and 75″. The photometric properties of our sample are consistent with satellites of MW-mass galaxies and dwarf galaxies within 3 Mpc (Pace24).
For our hosts, we identify fewer satellites per magnitude bin per host compared to MW-mass hosts and we find fewer candidates per magnitude bin for the SMC-mass hosts compared to the LMC-mass hosts. This agrees with CDM simulations that predict as the host’s halo mass increases, the stellar mass increases, and so to does the number of satellites (e.g. Dooley17; Dooley17a; Santos-Santos22). Our low-mass hosts also exhibit the trend observed among MW-mass hosts, where satellite abundance correlates with host stellar mass, extending this relationship into the dwarf host galaxy mass range.
From our candidate sample, we establish upper and lower estimates of the LF for low-mass galaxies. To determine the lower estimate, we analyze the central concentration of candidates around their hosts, estimating a rough lower bound of 1.60.7 (1.20.4 ) satellites per LMC-mass (SMC-mass) host with M9. The high-likelihood sample serves as the upper estimate, with 4.01.4 (2.10.6 ) candidates per LMC-mass (SMC-mass) host. Our upper and lower estimates bracket the predicted range for satellites of dwarf galaxies from Dooley17; Dooley17a.